OpenStudy (anonymous):

Will give medal and will fan I you help! James left a bin outside in his garden to collect rainwater. He notices that 1/4 gallon of water fills 2/3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin. Show your work.

OpenStudy (anonymous):

@texaschic101 @iGreen

OpenStudy (igreen):

1/4/2/3 = x/1

OpenStudy (igreen):

That's a proportion..

OpenStudy (igreen):

\(\dfrac{0.25}{0.\overline 6} = \dfrac{x}{1}\)

OpenStudy (texaschic101):

it might be less confusing if you put the parenthesis (1/4) / (2/3) = x/1 now cross multiply and solve for x

OpenStudy (anonymous):

ok so it would be 2/12

OpenStudy (igreen):

No..you multiplied wrong..

OpenStudy (igreen):

We can just multiply 1 to both sides.. (1/4 * 1) / (2/3) = x

OpenStudy (texaschic101):

(1/4) / (2/3) = x/1 cross multiply 2/3x = 1/4 x = (1/4) / (2/3) x = 1/4 * 3/2 x = ?? can you finish ?

OpenStudy (igreen):

Lol

OpenStudy (igreen):

An easier way is to just multiply 1 to both sides :P (1/4 * 1) / (2/3) = x (1/4) / (2/3) = x 1/4 * 3/2 = x

OpenStudy (anonymous):

yes so it would be 3/8

OpenStudy (texaschic101):

BINGO :)

OpenStudy (anonymous):

thanks guys you rock! so in words it would be (1/4) divided by (2/3)= x/1 then you have to multiply 1 to both sides which will equal (1/4) / (2/3) = x then we cross multiply 1/4 * 3/2 which equals 3/8

OpenStudy (igreen):

Yep, you got it.